OpenStudy (anonymous):

What is the inverse of ƒ? ƒ(x) = 8x

geerky42 (geerky42):

y = 8x Switch x and y to each other. x = 8y Now isolate y. y = x/8 x/8 is the inverse of f.

OpenStudy (anonymous):

@geerky42 so the answer is ƒ-1 (x) = 8x ?

geerky42 (geerky42):

x/8

geerky42 (geerky42):

One can find the inverse of function by switch x and y to each other then isolate y.

OpenStudy (anonymous):

so is it A,B,C, OR D

geerky42 (geerky42):

How should I know?

OpenStudy (anonymous):

A) ƒ-1 (x) = 8x B) ƒ-1 (x) = x + 8 C) ƒ-1 (x) = x - 8 D) ƒ-1 (x) = x 8

geerky42 (geerky42):

\[f^{-1}(x) = \dfrac{x}{8}\]It's D.